Define TELive) T(X.7,2)=(2x, 44452, 4y +32) by c) Diagonalize MCT) d) Find adj (MCT)) e) Find inverse of MCT)
Answer #1

Ques:- define TELCU,U) by T(X, Y, Z) = (2x, 44+52 AY +32) (c) digonalize MCT) (d) find adj (MCTÍ (e) find inverse of MCT) TE2 2 ОО 0 us 0 43 0 0 0 0 1 0 2-1 0 0 u-l 5 4 37 (2-1) [(4-1) (3 +) -20] +0101+0[0] = = (2-1) [12-72498-20) = 0 ► 2-1) (98-7)solue (1) lii) & (i) weget Lyco Uy + 4z U = - 43 Moto Eigenvertor corresponding to Arzt is u = 42 Now we find Eigen veetor CoNow eigen vector Corresponding to Pg=& (M-AI) W zo رقا 2 0 0 o us 0 43 olo 0 9 o-4 - 6W, = 0 (wis -4 Wy + 5 W₂=0 --(UV) 4 W₂1P1 - OL u matrix Transpose + [-5-4)= -9 Adj p = Co-factor matrx off we find Co-factor matrix for p a Now 1+1 = 0 11 = o u azNow Co-factor matrix of p Cuz aiz al A21 031 A23 A32 A33 0 0 -9 0 5 0 T 0 Adj p = 0 -4 O 0 0 کیا ؟ ou 5 0 Mow 0 -u pta 5 O 0Now Diagonal matrix of M (T) = 919 01-9 % -9 01-9 -10 1-9 0-9 Lol-g of g -71-91 Diagonal matrix D of M(T) = O 0 0 2. 0 0 & QuC23 = (-1) 243 an al 2 2 0 5 A31 a 32 0 u 0 012 Qu3 A22 923 1 0 u 5 C39 = (-0341 (22 = (-1) 3+2 jan 013 2 O 1 -10 1921 A23 0Now JM1 = 2 0 0 0 4 5 4 3 IM 1 = -16 Invesse MLT) In Adjm= - 11 -doo -10 0 -8 8 01-16 0/16 0/16 01-16 61-16 -101-16 01-16 -81